# Ikeda type lift on ${\rm SO}(3,n+1)$

### v1

@@ 4.2. Archimedean computation
For $n$ odd, we need some modification. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the discriminant of $\mathbb{Q}(\sqrt{q(\eta)})/\mathbb{Q}$. Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}(\sqrt{\pm q(\eta)})/\mathbb{Q}$. We also need the following the functional equation of Dirichlet $L$-function:
@@ 7. Fourier-Jacobi expansions of $E_{l}(g,\Phi_{\mathbb{f}})$ along $J$: an explicit form
Recall the Eisenstein series $E_{l}(g,\Phi_{f}):=E_{l}(g,\Phi_{\mathbb{f}},l+1)$ defined by the Pollack’s section $f_{l}(\ast,\Phi_{\mathbb{f}},l+1)\in I(\mathbf{1},l-\frac{n}{2})={\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}|\nu|^{l-\frac{n}{2}}$ (see Section 4). In this section, we explicitly compute the Fourier-Jacobi coefficients of $E_{l}(g,\Phi_{\mathbb{f}})$.
@@ 7.2. Computation of Fourier-Jacobi coefficients
Now, we are ready to compute the Fourier-Jacobi expansion of $E_{l}$ when $s=l+1$. Recall Section 5 and (4.1). Let $w_{1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&I_{n}&0\\ 1&0&0\end{array}\right)\in{\rm SO}(V^{\prime})(\mathbb{Q})$ so that
@@ 8.1.2. The odd case
$A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}}):=c(\mathfrak{d}_{\eta})\mathfrak{d}_{\eta}^{-\frac{k-1}{2}}q(\eta)^{\frac{k-1}{4}}\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}].$
@@ 8.3. Automorphy of the formal series
We give a proof only for $n$ even and the odd case is similarly proved. Let $f(z)=\sum_{n=1}^{\infty}a_{f}(n)n^{\frac{k-1}{2}}q^{n}$ be the newform in Theorem 1.1. For each prime $p$, let $a_{f}(p)=\alpha_{p}+\alpha^{-1}_{p}$. We will prove that the formal series
@@ 9. $L$-functions of Ikeda type lift
Let $f=\displaystyle\sum_{n=1}^{\infty}a_{f}(n)n^{\frac{k-1}{2}}q^{n}$ be a Hecke eigen cusp form of weight $k=\left\{\begin{array}[]{ll}l-\frac{n}{2}+1&\text{for $n$ even,}\\ 2l-n+1&\text{for $n$ odd}\end{array}\right$. with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $L(s,\pi_{f})=\displaystyle\prod_{p}L(s,\pi_{p})$, where $L(s,\pi_{p})=(1-\alpha_{p}p^{-s})^{-1}(1-\alpha_{p}^{-1}p^{-s})^{-1}$.
@@ 9.1. $n$ even
Let $n=8a+2$ ($n>2$). Let $F=F_{f}$ be the Ikeda type lift of $f$ on ${\rm SO}(3,1+n)$, and let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F$.
@@ 12. Appendix
In this section, we explicitly compute $C(S,r)=C(r,\sqrt{\beta}^{-1}\lambda)$ in (7.11) as a generalization of [22, Theorem A.1]: Let
@@
Let $I_{v}=I_{v}(r,\mu,\lambda)$. Then by the change of variables, we see that $I_{-v}=I_{v}$ and $\overline{I_{v}}=I_{v}$. So $I_{v}$ is real. Also [22, Proposition A.2] says, for any $r$,
@@
Now we explain how to modify the proof of [22, Proposition A.2] to obtain $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}.$ Then together with the recurrence relation above, we obtain the theorem.
Let $s=\zeta t^{\frac{1}{2}}+\zeta^{-1}t^{-\frac{1}{2}}$, where $\zeta=e^{\sqrt{-1}\frac{\pi}{4}}$. Then $s^{2}=2+\sqrt{-1}(t-t^{-1})$ and $s\bar{s}=t+t^{-1}$. Let $\alpha=\mu+\lambda^{2}$. As in the proof of [22, Proposition A.2], we have

### v2

@@ 4.2. Archimedean computation
For $n$ odd, we need some modification. Write $q(\eta)=\mathfrak{d}_{\eta}\mathfrak{f}_{\eta}^{2}$ such that $\mathfrak{d}_{\eta}$ is the absolute discriminant of $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$ as in the introduction. Let $\chi_{\eta}$ be the primitive Dirichlet character corresponding to $\mathbb{Q}\left(\sqrt{\epsilon q(\eta)}\right)/\mathbb{Q}$. We also need the following the functional equation of Dirichlet $L$-function:
@@ 7. Fourier-Jacobi expansions of $E_{l}(g,\Phi_{\mathbb{f}})$ along $J$: an explicit form
Recall the Eisenstein series $E_{l}(g,\Phi_{f}):=E_{l}(g,\Phi_{\mathbb{f}},l+1)$ defined by the Pollack’s section $f_{l}(\ast,\Phi_{\mathbb{f}},l+1)\in I(\mathbf{1},l-\frac{n}{2})={\rm Ind}^{G(\mathbb{A})}_{P(\mathbb{A})}|\nu|^{l-\frac{n+2}{2}}$ (see Section 4). In this section, we explicitly compute the Fourier-Jacobi coefficients of $E_{l}(g,\Phi_{\mathbb{f}})$.
@@ 7.2. Computation of Fourier-Jacobi coefficients
Now, we are ready to compute the Fourier-Jacobi expansion of $E_{l}$ when $s=l+1$. Recall the settings in Section 5 and (4.1). Let $w_{1}=\left(\begin{array}[]{ccc}0&0&1\\ 0&I_{n}&0\\ 1&0&0\end{array}\right)\in{\rm SO}(V^{\prime})(\mathbb{Q})$ so that
@@ 8.1.2. The odd case
$A(\eta;\{X_{p}\}_{p})(g_{\mathbb{f}}):=c(\mathfrak{d}_{\eta})\mathfrak{f}_{\eta}^{\frac{k-1}{2}}\prod_{p}\widetilde{Q}_{\eta,g_{p}\cdot\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}].$
@@ 8.3. Automorphy of the formal series
Let $f=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$ be a normalized Hecke eigen cusp form of weight $k=\left\{\begin{array}[]{ll}l-\frac{n}{2}+1&\text{for $n$ even,}\\ 2l-n+1&\text{for $n$ odd}\end{array}\right$. with respect to ${\rm SL}_{2}(\mathbb{Z})$. For each prime $p$, let $a_{f}(p)=\alpha_{p}+\alpha^{-1}_{p}$. We will prove that the formal series
@@
First we check the non-vanishing of $F_{f}$:
Lemma 8.1.
There exists $\eta\in V^{\prime}(\mathbb{Q})$ with $q(\eta)>0$ such that $A(\eta;\{\alpha_{p}\}_{p})\neq 0$.
Proof.
When $n$ is even, it is easy to find $\eta$ with $q(\eta)=1$. Thus, we have $A(\eta;\{\alpha_{p}\}_{p})=1$ by definition.
Now assume $n$ is odd. Let $D$ be a square-free fundamental discriminant. Then for $\eta=(D,D,\overbrace{0,\ldots,0}^{n},\pm 1,\pm 1)$, $q(\eta)=\pm D$, where we choose a sign so that $\pm D>0$. Then by Theorem 4.3, Theorem 4.9, and (8.3), $A(\eta;\{\alpha_{p}\}_{p})=c(|D|)$. By Kohnen-Zagier [18, Theorem 1], $c(|D|)$ is proportional to the central value of $f\otimes\chi_{D}$. The claim follows from [22, p.654]. ∎
Henceforth, we assume that $n$ is even. The odd case is similarly proved.
@@
8.4. Hecke Eigenform Property
[…]
Lemma 8.6.
The Eisenstein series $E_{l}$ is a Hecke eigenform.
Proof.
We may prove it for $T=T_{h}\in\mathcal{H}_{\mathbb{f}}$. By definition, we have
[…]
Since the finite part of $f$ is a non-zero section of the one-dimensional space $\Big({\rm Ind}^{G(\mathbb{A}_{\mathbb{f}})}_{P(\mathbb{A}_{\mathbb{f}})}|\nu|_{\mathbb{f}}^{l-\frac{n+2}{2}}\Big)^{G(\widehat{\mathbb{Z}})}$ which also preserved under the action of $\mathcal{H}_{\mathbb{f}}$. Thus, there exists $c_{h}\in\mathbb{C}^{\times}$ such that $T_{h}f=c_{h}f$. The claim follows. ∎
Proposition 8.7.
The automorphic form $F_{f}$ is a Hecke eigenform.
Proof.
We may prove it for $T=T_{h},\ h\in P(\mathbb{Q})$ . We write $G(\widehat{\mathbb{Z}})hG(\widehat{\mathbb{Z}})=\displaystyle\coprod_{i=1}^{r}h_{i}G(\widehat{\mathbb{Z}}),\ h_{i}\in P(\mathbb{Q})$. It is easy to see that $|\nu(h)|_{\mathbb{f}}=|\nu(h_{i})|_{\mathbb{f}}$ for any $i$. Let $l^{\prime}>n+1$ be a sufficiently large integer. Assume $n$ is even. Then, we see that
$T_{h}E_{l^{\prime}}(g)=\sum_{i}E_{l^{\prime}}(g_{\infty},g_{\mathbb{f}}h_{i}\Phi_{\mathbb{f}},s=l^{\prime}+1)=C_{l^{\prime},n}\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}\sum_{i=1}^{r}q(\eta)^{\frac{l^{\prime}-\frac{n}{2}}{2}}\prod_{p<\infty}\widetilde{Q}_{\eta,(g_{\mathbb{f}}h_{i})\cdot\Phi_{p}}(p^{\frac{l^{\prime}-\frac{n}{2}}{2}})W_{2\pi\eta}(g_{\infty})$
where $\widetilde{Q}_{\eta,\Phi_{p}}=\widetilde{Q}_{\eta,p}$. By Lemma 4.1, we have
$T_{h}E_{l^{\prime}}(g)=C_{l^{\prime},n}|\nu(h)|_{\mathbb{f}}^{-\frac{n+2}{2}}\sum_{\eta\in V^{\prime}(\mathbb{Q})\atop q(\eta)>0}\sum_{i=1}^{r}q(\eta)^{\frac{l^{\prime}-\frac{n}{2}}{2}}\prod_{p<\infty}\widetilde{Q}_{h^{-1}_{i}\eta,g_{\mathbb{f}}\cdot\Phi_{p}}(p^{\frac{l^{\prime}-\frac{n}{2}}{2}})W_{2\pi\eta}(g_{\infty}).$
We choose $\eta_{0}\in V^{\prime}(\mathbb{Q})$ in Lemma 8.1. By Lemma 8.6, there exists a non-zero constant $c_{h}$ such that $T_{h}E_{l^{\prime}}(g)=c_{h}E_{l^{\prime}}(g)$. Evaluating $g=g_{\infty}$ (hence $g_{\mathbb{f}}=1$), we see that $c_{h}=\alpha_{h}(\{p^{\frac{l^{\prime}-\frac{n}{2}}{2}}\}_{p})$ where
$\alpha_{h}(X):=|\nu(h)|_{\mathbb{f}}^{-\frac{n+2}{2}}\sum_{i=1}^{r}\prod_{p<\infty}\widetilde{Q}_{h^{-1}_{i}\eta_{0},\Phi_{p}}(X_{p})\in\otimes^{\prime}_{p<\infty}\mathbb{C}[X_{p},X^{-1}_{p}],\ X=\{X_{p}\}_{p}.$
By using [10, Lemma 10.1] and the argument for [15, Proposition 10.1], we see that
$T_{h}F_{f}=\alpha_{h}(\{\alpha_{p}\}_{p})F_{f}.$
The odd case is similarly proved and we omit the proof. ∎
@@ 9. $L$-functions of Ikeda type lift
Let $f=\displaystyle\sum_{m=1}^{\infty}a_{f}(m)m^{\frac{k-1}{2}}q^{m}$ be a Hecke eigen cusp form of weight $k=\left\{\begin{array}[]{ll}l-\frac{n}{2}+1&\text{for $n$ even,}\\ 2l-n+1&\text{for $n$ odd}\end{array}\right$. with respect to ${\rm SL}_{2}(\mathbb{Z})$. Let $a_{f}(p)=\alpha_{p}+\alpha_{p}^{-1}$. Let $L(s,\pi_{f})=\displaystyle\prod_{p}L(s,\pi_{p})$, where $L(s,\pi_{p})=(1-\alpha_{p}p^{-s})^{-1}(1-\alpha_{p}^{-1}p^{-s})^{-1}$.
@@ 9.1. $n$ even
Let $n=8a+2$ ($n>2$). Let $F=F_{f}$ be the Ikeda type lift of $f$ on ${\rm SO}(3,n+1)$, and let $\Pi_{F}=\otimes^{\prime}\Pi_{p}$ be the cuspidal representation associated with $F$.
@@ 12. Appendix
In this section, we explicitly compute $C(S,r)=C(r,\sqrt{\beta}^{-1}\lambda)$ in (7.11) as a generalization of [24, Theorem A.1]: Let
@@
Let $I_{v}=I_{v}(r,\mu,\lambda)$. Then by the change of variables, we see that $I_{-v}=I_{v}$ and $\overline{I_{v}}=I_{v}$. So $I_{v}$ is real. Also [24, Proposition A.2] says, for any $r$,
@@
Now we explain how to modify the proof of [24, Proposition A.2] to obtain $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}.$ Then together with the recurrence relation above, we obtain the theorem.
Let $s=\zeta t^{\frac{1}{2}}+\zeta^{-1}t^{-\frac{1}{2}}$, where $\zeta=e^{\sqrt{-1}\frac{\pi}{4}}$. Then $s^{2}=2+\sqrt{-1}(t-t^{-1})$ and $s\bar{s}=t+t^{-1}$. Let $\alpha=\mu+\lambda^{2}$. As in the proof of [24, Proposition A.2], we have